OpenStudy (anonymous):

Solve the equation ((-i)/(x-iy)) = ((4+7i)/(5-3i)) for x and y, leaving your answers as rational numbers. I made the denominator of both real so ((-x)/(x^2+y^2)) = (47/34) and ((y)/(x^2 + y^2)) = (-1/34) however I don't know where to go from here. Any help would be appreciated.

OpenStudy (anonymous):

Forgot to mention I was comparing the real and imaginary parts for the second bit.

OpenStudy (anonymous):

I got -47/34 and -1/34....

OpenStudy (anonymous):

I have -x on the numerator, I just didn't simplify that when I typed it. How would I go about solving for x and y? I seem to go in circles. Thanks

OpenStudy (anonymous):

Ah, Ok we agree then... let me have a look.

OpenStudy (anonymous):

It's a non-calculator question. I probably should've mentioned that.

OpenStudy (anonymous):

x^2 + y^2 = 1 ?

OpenStudy (anonymous):

Unit circle?

OpenStudy (anonymous):

Where did you get that from?

OpenStudy (anonymous):

It's a standard equation for a circle with radius 1. Or parametrized, cos^theta + sin^2 theta = 1.

OpenStudy (anonymous):

How did you get to x^2 + y^2 = 1? I can't get out of having y's. So, I have x^2 + y^2 = -34/y

OpenStudy (anonymous):

-34y sorry.

OpenStudy (anonymous):

You can regard your complex numbers as living in the complex plane. I'm not sure what you mean by a derivation. If I take a point x,y on the unit circle in the first quadrant, pythagorus tells u that x is cos theta and y is sin theta and cos^2 + sin^2 is 1. U can also get it from de Moivre's theorem. But it is accepted as standard.

OpenStudy (anonymous):

So x is 47/34 and y is -1/34, u already found the answer.

OpenStudy (anonymous):

-47/34, sorry.

OpenStudy (anonymous):

If ((x)/(x^2+y^2)) = (-47/34) and ((y)/(x^2 + y^2)) = (-1/34) Then surely x and y themselves will have different values. As x = (-47/34)(x^2+y^2) and y = (-1/34)(x^2 + y^2) Plugging those into WolframAlpha I get x as -47/65 and y as = -1/65 I can't work how to solve the equations for x and y however.

OpenStudy (anonymous):

U want x and y real, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So yuo scroll down the page to the alternate form (x, y real).

OpenStudy (anonymous):

That hasn't solved for x and y, that has just stated that ((ix)/(x^2+y^2)) = (-47i/34) and ((y)/(x^2 + y^2)) = (-1/34).

OpenStudy (anonymous):

That is telling you (by comparing real and imaginary parts as u did) that the answer is an alternate form of your equation. The x^2 +y^2 are gone because they equal 1. Not sure about the i artifact, could be because of the way u have written the question with -i at the beginning.

OpenStudy (anonymous):

Notice u get the same result if swap i for -i

OpenStudy (anonymous):

Let me try another, simpler tack. If I write 2x/5 = 3y/5 = 1/5 do u agree that I can cancel the 5's

OpenStudy (anonymous):

Treating the compared real and imaginary parts (without the i's) as simultaneous equations and plugging them in gets http://www.wolframalpha.com/input/?i=x+%3D+%28-47%2F34%29%28x^2%2By^2%29%2C+y+%3D+%28-1%2F34%29%28x^2+%2B+y^2%29 Which is the value for x and y.

OpenStudy (anonymous):

Also, yes. It's multiplying through by 5.

OpenStudy (anonymous):

I don't know what else to tell u, I think u have the answer already.

OpenStudy (anonymous):

Leave the question here, see if someone else will chime in....

OpenStudy (anonymous):

Ok, thanks for the help anyway.

OpenStudy (anonymous):

Alright, I figured this out now....

OpenStudy (anonymous):

The thing goes awry in your method. Simply cross multiply and multiply out. Then compare your real and imaginary parts and u get 2 simultaneous equations in x and y.

OpenStudy (anonymous):

I have 2 simultaneous equations at the moment which I can't solve. I can't get out of being in x's and y's. Are they different doing it that way? Generally, it seems to be better to make the denominator real first.

OpenStudy (anonymous):

No, that's what causes all the problems (the x^2 + y^2). Just do what I said you will see--y =-1/65, I didn't calculate x.

OpenStudy (anonymous):

Alright, I'll try it. Thanks.

OpenStudy (anonymous):

Ok, I've got it now. Thank you. Wish I tried that earlier though.

OpenStudy (anonymous):

U and me both...:-) Still, good workout.

OpenStudy (anonymous):

Yep.